# Stackelberg vs. Nash in the Lottery Colonel Blotto Game

### v1

@@ 7 Conclusion and Future Work
We investigate whether, and to what extent, moving first provides an advantage to a player in the Lottery Colonel Blotto game. To address this, we reduce the search space for the follower’s best response strategies and conduct a comparison between the Stackelberg equilibrium and the Nash equilibrium. This comparison is enabled by several insightful constructions and the water-filling understanding of the follower’s best response strategy.
Future research could extend the findings of this study on the Stackelberg version of the Colonel Blotto game in several ways. Additionally, incorporating scenarios with asymmetric information, where players have unequal knowledge about each other’s budgets or battlefield values, can provide insights into real-world applications.
@@ Appendix A OMITTED PROOFS FROM SECTION 3
If the leader’s commitment is $\bm{x}_{a}=(x_{a1},\cdots,x_{an})$, the follower’s best response is $\bm{x}_{b}=BR(\bm{x}_{a})=(x_{b1},\cdots,x_{bn})$ in $\mathcal{G}$, and the leader’s commitment is $\bm{x}_{a}^{(1)}=(x_{a1},\cdots,x_{a(r-1)},x_{a(r+1)},\cdots,x_{an},\frac{x_{ar}}{t},\cdots,\frac{x_{ar}}{t})$ in $\mathcal{G}^{(1)}$, we have $\bm{x}_{b}^{(1)}=BR(\bm{x}_{a}^{(1)})=(x_{b1},\cdots,x_{b(r-1)},x_{b(r+1)},\cdots,x_{bn},\frac{x_{br}}{t},\cdots,\frac{x_{br}}{t})$. Furthermore, we have that $u_{a}(\bm{x}_{a}^{(1)},\bm{x}_{b}^{(1)})=u_{a}(\bm{x}_{a},\bm{x}_{b})$.
@@
$\frac{x_{aj^{\prime}}^{(2)}v_{aj}^{(2)}}{x_{aj^{\prime}}^{(2)}+x_{bj^{\prime}}^{(2)}}+\frac{x_{aj}^{(2)}v_{aj^{\prime}}^{(2)}}{x_{aj}^{(2)}+x_{bj}^{(2)}}-\frac{x_{aj}^{(2)}v_{aj}^{(2)}}{x_{aj}^{(2)}+x_{bj}^{(2)}}-\frac{x_{aj^{\prime}}^{(2)}v_{aj^{\prime}}^{(2)}}{x_{aj^{\prime}}^{(2)}+x_{bj^{\prime}}^{(2)}}$ (15)
$=$ $(v_{aj^{\prime}}^{(2)}-v_{aj}^{(2)})\left(\frac{x_{aj}^{(2)}}{x_{aj}^{(2)}+x_{bj}^{(2)}}-\frac{x_{aj^{\prime}}^{(2)}}{x_{aj^{\prime}}^{(2)}+x_{bj^{\prime}}^{(2)}}\right).$
@@
$\begin{cases}x_{bj}^{(2)}=\frac{(x_{aj}^{(2)})^{\frac{1}{2}}}{\sum_{l\in K(\bm{x}_{a}^{(2)})}(x_{al}^{(2)})^{\frac{1}{2}}}(x_{b}+\sum_{l\in K(\bm{x}_{a}^{(2)})}x_{al}^{(2)})-x_{aj}^{(2)};\\ x_{bj^{\prime}}^{(2)}=\frac{(x_{aj^{\prime}}^{(2)})^{\frac{1}{2}}}{\sum_{l\in K(\bm{x}_{a}^{(2)})}(x_{al}^{(2)})^{\frac{1}{2}}}(x_{b}+\sum_{l\in K(\bm{x}_{a}^{(2)})}x_{al}^{(2)})-x_{aj^{\prime}}^{(2)}.\\ \end{cases}$
We substitute $x_{bj}^{(2)}$ and $x_{bj^{\prime}}^{(2)}$ into Equation (15), and we find that the value of Equation (15) is positive. Therefore, $u_{a}(\bm{x}_{a}^{(2)},\bm{x}_{b}^{(2)})<u_{a}(\bm{\hat{x}}_{a}^{(2)},\bm{\hat{x}}_{b}^{(2)})$. This immediately implies that for the leader, allocating at least the same amount of resources to the battlefields with higher value within his optimal commitment. ∎
@@ Appendix B OMITTED PROOFS FROM SECTION 4
Let us analyze the first subcase of Case 2, i.e., Case 2.1. Because of $K=[n]$, we have $\frac{1}{\beta^{2}}\sum_{j\in[n]}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}=x_{a}$. It follows that $\beta^{2}=\frac{\sum_{j\in[n]}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}{x_{a}}$. According to Problem (11), we have
$\max_{\bm{x}_{a}\in\bm{X}_{a}}\quad$ $\hat{u}_{a}(\bm{x}_{a})=\frac{\left(\sum_{j\in[n]}(\frac{v_{aj}}{v_{bj}}-\alpha)v_{aj}\right)\left(\sum_{j\in[n]}(v_{aj}-\alpha v_{bj})\right)}{\frac{\sum_{j\in[n]}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}{x_{a}}x_{b}+\sum_{j\in[n]}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}$ (21)
@@
$x_{aj}=(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}\cdot\frac{x_{a}}{\sum_{j\in[n]}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}},\;\forall j\in[n].$ (22)
To find a suitable parameter $\alpha$, we rewrite the Problem 21 as follows
$\max_{\bm{x}_{a}\in\bm{X}_{a}}\hat{u}_{a}(\bm{x}_{a})=\frac{\left((\sum_{j\in[n]}v_{aj})^{2}-(\sum_{j\in[n]}v_{bj})(\sum_{j\in[n]}\frac{v_{aj}^{2}}{v_{bj}})\right)\alpha}{\alpha^{2}(\sum_{j\in[n]}v_{bj})-2\alpha(\sum_{j\in[n]}v_{aj})+\sum_{j\in[n]}\frac{v_{aj}^{2}}{v_{bj}}}$ (23)
Note that $(\sum_{j\in[n]}v_{aj})^{2}-(\sum_{j\in[n]}v_{bj})(\sum_{j\in[n]}\frac{v_{aj}^{2}}{v_{bj}})<0$, we choose the $\alpha$ with negative value to maximize the leader’s utility. Solve the Problem 23, we have $\alpha=-\sqrt{\frac{1}{\sum_{j\in[n]}v_{bj}}\sum_{j\in[n]}\frac{v_{aj}^{2}}{v_{bj}}}$. Substitute $\alpha$ into Equation (22), the closed-form solution is obtained, that is
@@
Finally, let us analyse the sub-case 2.1. Because of $K\subsetneq[n]$, we can derive following equation
@@
Let $y=\beta^{2}$, we have $x_{aj}^{\frac{1}{2}}=(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha v_{bj}^{\frac{1}{2}})\frac{1}{\beta}=(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha v_{bj}^{\frac{1}{2}})\frac{1}{\sqrt{y}}$. Substituting $x_{aj}^{\frac{1}{2}}=(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha v_{bj}^{\frac{1}{2}})\frac{1}{\sqrt{y}}$ into the objective function of Problem 11, we obtain the leader’s utility, that is
$\hat{u}_{a}(\bm{x}_{a})$ $=\left(\sum_{j\in K}(\frac{v_{aj}}{v_{bj}}-\alpha)\frac{1}{\sqrt{y}}v_{aj}\right)\frac{\sum_{j\in K}(v_{aj}-\alpha v_{bj})\frac{1}{\sqrt{y}}}{\frac{1}{y}\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha v_{bj}^{\frac{1}{2}})^{2}+x_{b}}$
@@
$=\frac{\left(\sum_{j\in K}(\frac{v_{aj}}{v_{bj}}-\alpha)v_{aj}\right)\left(\sum_{j\in K}(v_{aj}-\alpha v_{bj})\right)}{yx_{b}+\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}$ (25)
@@
$\hat{y}=\frac{B_{1}\alpha^{2}+B_{2}\alpha+B_{3}-(v_{aK}-v_{bK}\alpha)\sqrt{B_{4}\alpha^{2}+B_{5}\alpha+B_{6}}}{2x_{b}^{2}v_{b\overline{K}}}$
When $B_{4}\alpha^{2}+B_{5}\alpha+B_{6}\geq 0$, $y$ has real number solutions. Substitute $\hat{y}$ into Equation (25), and obtain the following optimization problem
$\max_{\bm{x_{a}}\in\bm{X}_{a}}\quad$ $\frac{\left(\sum_{j\in K}(\frac{v_{aj}}{v_{bj}}-\alpha)v_{aj}\right)\left(\sum_{j\in K}(v_{aj}-\alpha v_{bj})\right)}{\hat{y}x_{b}+\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}$ (27)
@@
We need to find a suitable parameter $\alpha$ that maximizes the leader’s utility. Therefore, the optimal commitment is as follows
$x_{aj}=\begin{cases}\frac{2x_{b}^{2}v_{b\overline{K}}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}{B_{1}\alpha^{2}+B_{2}\alpha+B_{3}-(v_{aK}-v_{bK}\alpha)\sqrt{B_{4}\alpha^{2}+B_{5}\alpha+B_{6}}},\;j\in K;\\ \frac{v_{bj}\left(x_{b}+\sum_{l\in K}x_{al}\right)^{2}}{\left(\sum_{l\in K}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}},\;j\in\overline{K}.\end{cases}$
@@
$u_{a}(\hat{\bm{x}}_{a}(\varepsilon))-u_{a}(\bm{x}_{a})=$ $\frac{\varepsilon}{x_{aK}+x_{b}}\Bigg{[}-\sum_{j\in K}(\frac{x_{aj}}{v_{bj}})^{\frac{1}{2}}\sum_{j\in K}(x_{aj}v_{bj})^{\frac{1}{2}}+x_{aK}v_{an}+(\frac{1-\varepsilon}{\varepsilon})^{\frac{1}{2}}\Big{[}(x_{aK}v_{bn})^{\frac{1}{2}}\sum_{j\in K}(\frac{x_{aj}}{v_{bj}})^{\frac{1}{2}}+$
$(\frac{x_{aK}}{v_{bn}})^{\frac{1}{2}}v_{an}\sum_{j\in K}(x_{aj}v_{bj})^{\frac{1}{2}}\Big{]}\Bigg{]}$

### v3

@@ 7 Conclusion and Future Work
We explore the advantages of the first mover in the Lottery Colonel Blotto game. To address this, we reduce the number of supports in the follower’s best response strategies from $12^{n}-$ 1 to $n$. We derive a method to calculate the leader’s commitment for each possible support, allowing selection of the optimal one from these $n$ potential commitments. In addition, we provide the necessary and sufficient conditions under which the Stackelberg equilibrium and Nash equilibrium are equivalent. We find that the two equilibria are equivalent when the budget ratio between the leader and the follower is exactly equal to a certain functional value, which depends on the ratio of the valuations of the battlefields by the leader and the follower. Furthermore, we analyze the ratio of the leader’s utility under the Stackelberg equilibrium to that under the Nash equilibrium.
This work opens several directions for further research. First, analyzing the ratio between the follower’s utility under Stackelberg and Nash equilibria may yield deeper understanding of the incentives and trade-offs involved. Second, it is of interest to study which player benefits more from committing early, and under what conditions a player prefers to act as the leader or the follower. Finally, extending the model to settings with multiple followers introduces new layers of strategic complexity, where the interaction among followers becomes both theoretically rich and analytically challenging.
Acknowledgments
We thank the fnancial support from National Natural Science Foundation of China (Grant No. 62172422 and 72192805), and Key Laboratory of Interdisciplinary Research of Computation and Economics (Shanghai University of Finance and Economics), Ministry of Education. Jie Zhang was partially supported by a Leverhulme Trust Research Project Grant (2021 – 2024) and the EPSRC grant (EP/W014912/1).
@@ Appendix A OMITTED PROOFS FROM SECTION 3
If the leader’s commitment is $\bm{x}_{a}=(x_{a1},\cdots,$ $x_{an})$, the follower’s best response is $\bm{x}_{b}=BR(\bm{x}_{a})=(x_{b1},\cdots,x_{bn})$ in $\mathcal{G}$, and the leader’s commitment is $\bm{x}_{a}^{(1)}=(x_{a1},\cdots,x_{a(r-1)},x_{a(r+1)},\cdots,x_{an},\frac{x_{ar}}{t},\cdots,\frac{x_{ar}}{t})$ in $\mathcal{G}^{(1)}$, we have $,\bm{x}_{b}^{(1)}=BR(\bm{x}_{a}^{(1)})=(x_{b1},\cdots,x_{b(r-1)},x_{b(r+1)},$$\cdots,x_{bn},\frac{x_{br}}{t},\cdots,\frac{x_{br}}{t})$. Furthermore, we have that $u_{a}(\bm{x}_{a}^{(1)},$ $\bm{x}_{b}^{(1)})=u_{a}(\bm{x}_{a},\bm{x}_{b})$.
@@
$\frac{x_{aj^{\prime}}^{(2)}v_{aj}^{(2)}}{x_{aj^{\prime}}^{(2)}+x_{bj^{\prime}}^{(2)}}+\frac{x_{aj}^{(2)}v_{aj^{\prime}}^{(2)}}{x_{aj}^{(2)}+x_{bj}^{(2)}}-\frac{x_{aj}^{(2)}v_{aj}^{(2)}}{x_{aj}^{(2)}+x_{bj}^{(2)}}-\frac{x_{aj^{\prime}}^{(2)}v_{aj^{\prime}}^{(2)}}{x_{aj^{\prime}}^{(2)}+x_{bj^{\prime}}^{(2)}}$
$=$ $(v_{aj^{\prime}}^{(2)}-v_{aj}^{(2)})\left(\frac{x_{aj}^{(2)}}{x_{aj}^{(2)}+x_{bj}^{(2)}}-\frac{x_{aj^{\prime}}^{(2)}}{x_{aj^{\prime}}^{(2)}+x_{bj^{\prime}}^{(2)}}\right).$ (15)
@@
$\begin{cases}x_{bj}^{(2)}=\frac{(x_{aj}^{(2)})^{\frac{1}{2}}}{\sum_{l\in K(\bm{x}_{a}^{(2)})}(x_{al}^{(2)})^{\frac{1}{2}}}(x_{b}+\sum_{l\in K(\bm{x}_{a}^{(2)})}x_{al}^{(2)})-x_{aj}^{(2)};\\ x_{bj^{\prime}}^{(2)}=\frac{(x_{aj^{\prime}}^{(2)})^{\frac{1}{2}}}{\sum_{l\in K(\bm{x}_{a}^{(2)})}(x_{al}^{(2)})^{\frac{1}{2}}}(x_{b}+\sum_{l\in K(\bm{x}_{a}^{(2)})}x_{al}^{(2)})-x_{aj^{\prime}}^{(2)}.\end{cases}$
We substitute $x_{bj}^{(2)}$ and $x_{bj^{\prime}}^{(2)}$ into Equation (A), and we find that the value of Equation (A) is positive. Therefore, $u_{a}(\bm{x}_{a}^{(2)},\bm{x}_{b}^{(2)})<u_{a}(\bm{\hat{x}}_{a}^{(2)},\bm{\hat{x}}_{b}^{(2)})$. This immediately implies that for the leader, allocating at least the same amount of resources to the battlefields with higher value within his optimal commitment. ∎
@@ Appendix B OMITTED PROOFS FROM SECTION 4
For the sake of simplicity, we define the following expressions.
Definition 1.
For $i\in\{a,b\}$, $K\subseteq[n]$, let
$v_{iK}$ $=\sum_{j\in K}v_{ij},$
$c_{K}$ $=\sum_{j\in K}\frac{v_{aj}^{2}}{v_{bj}},$
$B_{1}(K)$ $=x_{a}v_{bK}^{2}-2x_{b}v_{bK}v_{b\overline{K}},$
$B_{2}(K)$ $=4x_{b}v_{aK}v_{b\overline{K}}-2x_{a}v_{aK}v_{bK},$
$B_{3}(K)$ $=x_{a}v_{aK}^{2}-2x_{b}v_{b\overline{K}}c_{K},$
$B_{4}(K)$ $=x_{a}^{2}v_{bK}^{2}-4x_{b}(x_{a}+x_{b})v_{bK}v_{b\overline{K}},$
$B_{5}(K)$ $=8x_{b}(x_{a}+x_{b})v_{aK}v_{b\overline{K}}-2x_{a}^{2}v_{aK}v_{bK},$
$B_{6}(K)$ $=x_{a}^{2}v_{aK}^{2}-4x_{b}(x_{a}+x_{b})c_{K}v_{b\overline{K}},$
$\phi_{1}(\theta,K)$ $=B_{1}(K)\theta^{2}+B_{2}(K)\theta+B_{3}(K),$
$\phi_{2}(\theta,K)$ $=B_{4}(K)\theta^{2}+B_{5}(K)\theta+B_{6}(K).$
We consider three cases, namely
  • Case 1. $\forall j,h\in K$, $\frac{v_{aj}}{v_{bj}}=\frac{v_{ah}}{v_{bh}}$, or $K=\{1\}$.
  • Case 2.1. there exist $j,h\in K$, such that $\frac{v_{aj}}{v_{bj}}\neq\frac{v_{ah}}{v_{bh}}$, and $K=[n]$.
  • Case 2.2. there exist $j,h\in K$, such that $\frac{v_{aj}}{v_{bj}}\neq\frac{v_{ah}}{v_{bh}}$, and $K\subsetneq[n]$.
Rearranging Equation (21), we obtain
This can be calculated in $O(n)$ time.
Second, we consider the case where $\frac{v_{aj}}{v_{bj}}\neq\frac{v_{ah}}{v_{bh}}$, leading to cases 2.1 and 2.2. In this case, $\beta\neq 0$. According to Lemma 3, for $\forall j\in K$, we have $(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha v_{bj}^{\frac{1}{2}})\frac{1}{\beta}=x_{aj}^{\frac{1}{2}}$. The resources invested by the leader in $K$ are $\frac{1}{\beta^{2}}\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}$. For $\forall j\in\overline{K}$, we have
Let us analyze Case 2.1. Because of $K=[n]$, we have $\frac{1}{\beta^{2}}\sum_{j\in[n]}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}=x_{a}$. It follows that $\beta^{2}=\frac{\sum_{j\in[n]}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}{x_{a}}$. According to Problem (12), we have
$\max_{\bm{x}_{a}\in\bm{X}_{a}}$ $\hat{u}_{a}(\bm{x}_{a})=\frac{\left(\sum_{j\in[n]}(\frac{v_{aj}}{v_{bj}}-\alpha)v_{aj}\right)\left(\sum_{j\in[n]}(v_{aj}-\alpha v_{bj})\right)}{\frac{\sum_{j\in[n]}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}{x_{a}}x_{b}+\sum_{j\in[n]}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}$ (22)
@@
$x_{aj}=\frac{(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}\cdot x_{a}}{\sum\limits_{j\in[n]}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}},\;\forall j\in[n].$ (23)
To find a suitable parameter $\alpha$, we rewrite the Problem 22 as follows
$\max\limits_{\bm{x}_{a}\in\bm{X}_{a}}\hat{u}_{a}(\bm{x}_{a})=\frac{\left((\sum\limits_{j\in[n]}v_{aj})^{2}-(\sum\limits_{j\in[n]}v_{bj})(\sum\limits_{j\in[n]}\frac{v_{aj}^{2}}{v_{bj}})\right)\alpha}{\alpha^{2}(\sum\limits_{j\in[n]}v_{bj})-2\alpha(\sum\limits_{j\in[n]}v_{aj})+\sum\limits_{j\in[n]}\frac{v_{aj}^{2}}{v_{bj}}}$ (24)
Note that $(\sum_{j\in[n]}v_{aj})^{2}-(\sum_{j\in[n]}v_{bj})(\sum_{j\in[n]}\frac{v_{aj}^{2}}{v_{bj}})<0$, we choose the $\alpha$ with negative value to maximize the leader’s utility. Solve the Problem 24, we have $\alpha=-\sqrt{\frac{1}{\sum_{j\in[n]}v_{bj}}\sum_{j\in[n]}\frac{v_{aj}^{2}}{v_{bj}}}$. Substitute $\alpha$ into Equation (23), the closed-form solution is obtained, that is
@@
This can be calculated in $O(n)$ time.
Finally, let us analyse Case 2.2. Because of $K\subsetneq[n]$, we can derive following equation
@@
Let $y=\beta^{2}$, we have $x_{aj}^{\frac{1}{2}}=(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha v_{bj}^{\frac{1}{2}})\frac{1}{\beta}=(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha v_{bj}^{\frac{1}{2}})\frac{1}{\sqrt{y}}$. Substituting $x_{aj}^{\frac{1}{2}}=(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha v_{bj}^{\frac{1}{2}})\frac{1}{\sqrt{y}}$ into the objective function of Problem 12, we obtain the leader’s utility, that is
$\hat{u}_{a}(\bm{x}_{a})$ $=\left(\sum_{j\in K}(\frac{v_{aj}}{v_{bj}}-\alpha)\frac{1}{\sqrt{y}}v_{aj}\right)\frac{\sum_{j\in K}(v_{aj}-\alpha v_{bj})\frac{1}{\sqrt{y}}}{\frac{1}{y}\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha v_{bj}^{\frac{1}{2}})^{2}+x_{b}}$ (26)
@@
$=\frac{\left(\sum_{j\in K}(\frac{v_{aj}}{v_{bj}}-\alpha)v_{aj}\right)\left(\sum_{j\in K}(v_{aj}-\alpha v_{bj})\right)}{yx_{b}+\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}$
@@
$\hat{y}=\frac{\phi_{1}(\alpha,K)-(v_{aK}-v_{bK}\alpha)\sqrt{\phi_{2}(\alpha,K)}}{2x_{b}^{2}v_{b\overline{K}}}$
When $\phi_{2}(\alpha,K)\geq 0$, $y$ has real number solutions. Substitute $\hat{y}$ into Equation (26), and obtain the following optimization problem
$\max_{\bm{x_{a}}\in\bm{X}_{a}}$ $\hat{u}(\bm{x}_{a})=\frac{\left(\sum_{j\in K}(\frac{v_{aj}}{v_{bj}}-\alpha)v_{aj}\right)\left(\sum_{j\in K}(v_{aj}-\alpha v_{bj})\right)}{\hat{y}x_{b}+\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}$ (29)
@@
$x_{aj}=\begin{cases}\frac{2x_{b}^{2}v_{b\overline{K}}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}{\phi_{1}(\alpha,K)-(v_{aK}-v_{bK}\alpha)\sqrt{\phi_{2}(\alpha,K)}},\;j\in K;\\ \frac{v_{bj}\left(x_{b}+\sum_{l\in K}x_{al}\right)^{2}}{\left(\sum_{l\in K}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}},\;j\in\overline{K}.\end{cases}$
This can be calculated in $O(n)$ time. ∎
@@
$u_{a}(\hat{\bm{x}}_{a}(\varepsilon))-u_{a}(\bm{x}_{a})$
$=$ $\frac{\varepsilon}{x_{aK}+x_{b}}\Bigg{[}-\sum_{j\in K}(\frac{x_{aj}}{v_{bj}})^{\frac{1}{2}}\sum_{j\in K}(x_{aj}v_{bj})^{\frac{1}{2}}+x_{aK}v_{an}+(\frac{1-\varepsilon}{\varepsilon})^{\frac{1}{2}}\Big{[}(x_{aK}v_{bn})^{\frac{1}{2}}\sum_{j\in K}(\frac{x_{aj}}{v_{bj}})^{\frac{1}{2}}+(\frac{x_{aK}}{v_{bn}})^{\frac{1}{2}}v_{an}\sum_{j\in K}(x_{aj}v_{bj})^{\frac{1}{2}}\Big{]}\Bigg{]}$